John Rider wants to accumulate $80,000 to be used for his daughter’s college education. He would like to have the amount available on December 31, 2026. Assume that the funds will accumulate in a certificate of deposit paying 8% interest compounded annually. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) Answer each of the following independent questions.

Required:

If John were to deposit a single amount, how much would he have to invest on December 31, 2021?

If John were to make five equal deposits on each December 31, beginning a year later, on December 31, 2022, what is the required amount of each deposit?

If John were to make five equal deposits on each December 31, beginning now, on December 31, 2021, what is the required amount of each deposit? (For all requirements, Round your final answers to nearest whole dollar amount.)
Answer #1

Single amount required to be deposited = Future value*Present value factor

= 80,000*PVF(8%, 5 years)

= 80,000*0.6806

= $54,448

Annual deposit required = Future value/Future value annuity factor

= 80,000/5.8666

= $13,636.52

Required amount of each deposit = Future amount/Future value of annuity due factor

= 80,000/6.3359

= $12,626.46